# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 8. Infinite sums of finitely generated projective modules
In this section we keep working with an algebra $\Lambda$ over an $h$-local domain $R$, with field of fractions $Q$ and such that $\Lambda_{Q}$ is simple artinian.
We keep using the notation of the previous section. In particular, the one introduced in Notation 7.3. As in Proposition 7.6, let $\mathcal{A}=\Phi^{\prime}(V(\Lambda))$. If $\ell$ is either 00 or $1$, then every projective module is a direct sum of finitely generated ones (cf. Remark 7.7). So we assume $\ell\geq 2$.
Let $\mathcal{B}$ be the image of $B(\Lambda)=V^{*}(\Lambda)\setminus V(\Lambda)$ in $\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$.
By Proposition 7.1, the genus classifies the modules of $B(\Lambda)$ up to isomorphism. Therefore, the modules in $B(\Lambda)$ that are isomorphic to a direct sum of finitely generated modules are the ones such that the image of its isomorphism class via $\Phi^{\prime}$ lies in
@@
In this section we will examine the question whether any projective $\Lambda$-module is a direct sum of finitely generated ones. By Proposition 2.8, we only need to characterize when finitely generated projective modules modulo an ideal of $\mathcal{T}(\Lambda)$ lift to finitely generated projective modules. Our next Lemma shows that we only need to ensure this property for the minimal elements in $\mathcal{T}(\Lambda)\setminus\{0\}$.
@@
    For any $I\in\mathcal{T}(\Lambda)$ and any finitely generated projective $\Lambda/I$-module $P^{\prime}$ there exists a finitely generated projective module $\Lambda$-module $P$ such that $P^{\prime}\cong P/PI$.
@@
    For any $I$ which is a minimal element of $\mathcal{T}(\Lambda)\setminus\{0\}$ and any finitely generated projective $\Lambda/I$-module $P^{\prime}$ there exists a finitely generated projective module $\Lambda$-module $P$ such that $P^{\prime}\cong P/PI$.
Proof.  Of course, only $(ii)\Rightarrow(i)$ needs a proof. To prove $(ii)\Rightarrow(i)$, let $I$ be a nonzero trace of $\Lambda$ containing a minimal trace $X\in T(\Lambda)\setminus\{0\}$. Such an $X$ exists by Corollary 6.10.
As $X_{Q}=\Lambda_{Q}$, $\Lambda/X$ is a torsion $R$-module. As $R$ is $h$-local, the canonical homomorphism $\Lambda/X\to\prod_{\mathfrak{m}\in\operatorname{mSpec}R}(\Lambda/X)_{\mathfrak{m}}$ induces an isomorphism $\varphi\colon\Lambda/X\to\bigoplus_{\mathfrak{m}\in\operatorname{mSpec}R}(\Lambda/X)_{\mathfrak{m}}$. Notice that, by Lemma 5.8, $X_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ for almost all maximal ideals $\mathfrak{m}$ of $R$.
Since $(\Lambda/X)_{\mathfrak{m}}\cong\Lambda_{\mathfrak{m}}/X_{\mathfrak{m}}$ is a semiperfect ring, the module $\bigoplus_{\mathfrak{m}\in\operatorname{mSpec}R}(\Lambda/X)_{\mathfrak{m}}$ can be seen as a finite product of semiperfect rings and $\varphi$ is an isomorphism of rings. In particular, $\Lambda/X$ is semiperfect.
@@
Next Lemma is going to be useful to be able to check when the criteria of Lemma 8.1 is satisfied.
@@
which has a solution in $\mathbb{N}$ if and only if the equation $my-sx_{\ell}=r_{\ell}$ has a solution in $\mathbb{N}$. Also, it is clear that the solutions are related as claimed in the statement.
@@
Under the Notation 7.3, the following are equivalent
@@
    For every choosing $a_{1},a_{2},\dots,a_{\ell}$, $1\leq a_{i}\leq t_{i}$, and for every $j\in\{1,\dots,\ell\}$ and $1\leq b\neq a_{j}\leq t_{j}$ there are non-negative integers $x_{1},x_{2},\dots,x_{\ell}$ such that $r_{j,b}+x_{j}r_{j,a_{j}}=x_{i}r_{i,a_{i}}$ for every $i\neq j$.
@@
    For every choosing $a_{1},a_{2},\dots,a_{\ell}$, $1\leq a_{i}\leq t_{i}$, and for every $j\in\{1,\dots,\ell\}$ and $1\leq b\neq a_{j}\leq t_{j}$ then $\gcd(\mathrm{lcm}(\{r_{1,a_{1}},\dots,r_{\ell,a_{\ell}}\}\setminus\{r_{j,a_{j}}\}),r_{j,a_{j}})$ divides $r_{j,b}$.
@@
Proof.  Assume $(i)$ is true, so finitely generated projective modules of $\Lambda$ modulo a (minimal) element of $\mathcal{T}(\Lambda)\setminus\{0\}$ can be lifted to finitely generated projective modules $\Lambda$-modules.
@@
By Remark 6.9, $\Lambda$ has an ideal $I\in\mathcal{T}(\Lambda)$ such that $I_{\mathfrak{m}_{i}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}_{i}}}(U_{i,a_{i}})$ for each $i=1,\dots,\ell$. As $\Lambda_{Q}$ is simple artinian and $R$ is $h$-local, $\Lambda/I\cong\bigoplus_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$. Therefore, every (finitely generated) projective $\Lambda/I$-module is of the form $\bigoplus_{i=1}^{\ell}M_{i}$, where $M_{i}$ is a (finitely generated) projective $(\Lambda/I)_{\mathfrak{m}_{i}}$-module. Note that $(\Lambda/I)_{\mathfrak{m}_{i}}\cong\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$ which is a semiperfect ring, so projective modules over it are described in Proposition 6.2 (iv).
@@
Consider a projective $\Lambda/I$-module $P^{\prime}$ such that $P_{\mathfrak{m}_{j}}^{\prime}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$ and $P_{\mathfrak{m}_{i}}^{\prime}=0$ if $i\neq j$. By assumption, there exists $P$ a finitely generated projective $\Lambda$-module such that $P/PI\cong P^{\prime}$. Then $P_{\mathfrak{m}_{i}}\cong U_{i,a_{i}}^{x_{i}}$ if $i\neq j$ and $P_{\mathfrak{m}_{j}}\cong U_{j,b}\oplus U_{j,a_{j}}^{x_{j}}$. Since $(P_{\mathfrak{m}_{i}})_{Q}\cong P_{Q}\cong(P_{\mathfrak{m}_{j}})_{Q}$, we get $x_{i}r_{i,a_{i}}=r_{j,b}+x_{j}r_{j,a_{j}}$. So $(ii)$ holds.
@@
Conversely, let $I$ be a minimal nonzero trace of $\Lambda$. Then $I_{\mathfrak{m}_{i}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}_{i}}}(U_{i,a_{i}})$ for some $a_{i}$, $i\in\{1,\dots,\ell_{i}\}$. Recall that $\Lambda/I\cong\bigoplus_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$ is a semiperfect ring, hence it is sufficient to lift an indecomposable projective $\Lambda/I$-module to a finitely generated projective $\Lambda$-module. So let $P^{\prime}$ be a finitely generated indecomposable projective $\Lambda/I$-module. Regarding the decomposition of $\Lambda/I$, $P^{\prime}_{\mathfrak{m}_{i}}=0$ for all $i$’s with exactly one exception. Assume $P_{\mathfrak{m}_{j}}^{\prime}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$, where $b\neq a_{j}$. Take $x_{1},\dots,x_{\ell}\in\mathbb{N}_{0}$ such that the equations in $(ii)$ are satisfied. Let $r=r_{j,b}+x_{j}r_{j,a_{j}}$.
@@
For $i=1,\dots,\ell$ set $U(\mathfrak{m}_{i})=U_{i,a_{i}}^{x_{i}}$ if $i\neq j$ and $U(\mathfrak{m}_{j})=U_{j,b}\oplus U_{j,a_{j}}^{x_{j}}$. For $\mathfrak{n}\in\operatorname{mSpec}(R)\setminus\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$ let $U(\mathfrak{n})$ be a projective $\Lambda_{\mathfrak{n}}$-module such that $U(\mathfrak{n})_{Q}$ has length $r$ (it exists by Lemma 6.8). Because of the choice of $x_{1},\dots,x_{\ell}$, we can use Proposition 7.2 to see the existence of a finitely generated projective $\Lambda$-module $P$ such that $P_{\mathfrak{m}}\cong U(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. Then $P_{\mathfrak{m}_{i}}I_{\mathfrak{m}_{i}}=P_{\mathfrak{m}_{i}}$ if $i\neq j$ and $P_{\mathfrak{m}_{j}}/P_{\mathfrak{m}_{j}}I_{\mathfrak{m}_{j}}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$. Therefore $P/PI\cong\bigoplus_{i=1}^{\ell}P_{\mathfrak{m}_{i}}/P_{\mathfrak{m}_{i}}I_{\mathfrak{m}_{i}}\cong P^{\prime}$. This proves $(ii)\Rightarrow(i)$.
@@
Under the assumptions in Notation 7.3, let $\ell=2$, and assume that $U_{1,1},U_{1,2}$ and $U_{2,1},U_{2,2}$ are representatives of the finitely generated indecomposable projective right modules over $\Lambda_{\mathfrak{m}_{1}}$ and $\Lambda_{\mathfrak{m}_{2}}$, respectively. Suppose that
@@
Notice that a locally semiperfect algebra $\Lambda$ exists by Example 7.4.
@@
Under the assumptions in Notation 7.3, there are examples of algebras $\Lambda$ where there is a projective $\Lambda$-module which is not a direct sum of finitely generated modules. Set now
@@
By Example 7.4, we know there exist locally semiperfect algebras $\Lambda$ with such invariants.
@@
Let $[P]\in B(\Lambda)$ be such that $\Phi^{\prime}([P])=((0,\infty),(1,\infty))$. Then there is no $\mathbf{X}\in\mathbb{N}_{0}^{2}\times\mathbb{N}_{0}^{2}$ of the form $\mathbf{X}=((0,x_{1,2}),(1,x_{2,2}))$ being a solution of (14) since $8x_{1,2}\equiv 0\mod 2$ and $1+4x_{2,2}\equiv 1\mod 2$. Therefore, there is no finitely generated projective $\Lambda$-module $P^{\prime}$ such that $\Phi^{\prime}([P^{\prime}])=((0,x_{1,2}),(1,x_{2,2}))$, so $P$ is a countably generated projective $\Lambda$-module which is not a direct sum of finitely generated modules.
Let $P_{1}$ be a finitely generated $\Lambda$-module such that $\Phi^{\prime}([P_{1}])=((0,1),(0,2))_{8}$, and let $P_{2}$ be such that $\Phi^{\prime}([P_{2}])=((0,1),(2,4))_{8}$. By Proposition 7.1, $P\cong P\oplus P_{1}^{(\omega)}$ and $P^{2}\cong P_{2}\oplus P_{1}^{(\omega)}$. Thus $P^{2}$ is a direct sum of finitely generated projective modules.

### v2

@@ 7. Rank and Genus of countably generated projective modules
Remark 7.10.
In Proposition 7.9, the condition on $\mathbf{Y}\in\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$ is equivalent to saying that, for each $i\in\{1,\dots,\ell\}$, at least one of the components of $(y_{i,1},\dots,y_{i,t_{i}})$ is $\infty$.
We have kept the notation of a system of linear equations because it is consistent with the characterization of the monoids that can be realized as $V(\Lambda)\sqcup B(\Lambda)$ for $\Lambda$ a semilocal ring given in [crelle]. Namely, the description of $V(\Lambda)\sqcup B(\Lambda)$ for $\Lambda$ a semilocal ring can always be given in terms of the solutions in ${\mathbb{N}}_{0}^{*}$ of a system of linear equations and linear congruences with coefficients in ${\mathbb{N}}_{0}$. However, it is not at all clear which invariants of the semilocal ring determine the coefficients of the system defining the monoid of relatively big projectives.
In our situation, by Proposition 7.2, $\Lambda_{\Sigma}$ where $\Sigma=R\setminus\bigcup_{i=1}^{\ell}\mathfrak{m}_{i}$ is a locally semiperfect, semilocal algebra over the semilocal domain $R_{\Sigma}$. We are proving that, at least for the genus, the system of equations that determines the monoid can be deduced from the structural properties of $\Lambda$ and its localizations at the maximal ideals of $R$. Note that if $\Lambda$ is finitely generated over $R$, then, by Proposition 7.2, $V^{*}(\Lambda_{\Sigma})=V(\Lambda)\sqcup B(\Lambda)\cong\Phi^{\prime}(V^{*}(\Lambda))$.
@@ 8. Infinite sums of finitely generated projective modules
In this section, we continue working with an algebra $\Lambda$ over an $h$-local domain $R$, with field of fractions $Q$, and such that $\Lambda_{Q}$ is simple artinian.
We keep using the notation from the previous section. In particular, the one introduced in Notation 7.4. As in Proposition 7.7, let $\mathcal{A}=\Phi^{\prime}(V(\Lambda))$. If $\ell$ is either 0 or $1$, then every projective module is a direct sum of finitely generated ones (cf. Remark 7.8). So we assume $\ell\geq 2$.
Let $\mathcal{B}$ be the image via $\Phi^{\prime}$ of $B(\Lambda)=V^{*}(\Lambda)\setminus V(\Lambda)$ in $\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$.
By Proposition 7.1, the genus classifies the modules of $B(\Lambda)$ up to isomorphism. Therefore, the modules in $B(\Lambda)$ that are isomorphic to a direct sum of finitely generated modules are those such that the image of their isomorphism class via $\Phi^{\prime}$ lies in
@@
In this section we will examine the question of whether any projective $\Lambda$-module is a direct sum of finitely generated modules. By Proposition 1.19 and Lemma 6.7, we only need to characterize when finitely generated projective modules modulo an ideal of $\mathcal{T}(\Lambda)$ lift to finitely generated projective modules. Our next Lemma shows that we only need to ensure this property for the minimal elements in $\mathcal{T}(\Lambda)\setminus\{0\}$.
@@
    For any $I\in\mathcal{T}(\Lambda)$ and any finitely generated projective $\Lambda/I$-module $P^{\prime}$, there exists a finitely generated projective $\Lambda$-module $P$ such that $P^{\prime}\cong P/PI$.
@@
    For any $I$ which is a minimal element of $\mathcal{T}(\Lambda)\setminus\{0\}$ and any finitely generated projective $\Lambda/I$-module $P^{\prime}$, there exists a finitely generated projective $\Lambda$-module $P$ such that $P^{\prime}\cong P/PI$.
Proof. Of course, only $(ii)\Rightarrow(i)$ needs a proof. To prove $(ii)\Rightarrow(i)$, let $I$ be a non-zero trace of $\Lambda$ containing a minimal trace $X\in T(\Lambda)\setminus\{0\}$. Such an $X$ exists by Corollary 6.11.
By Lemma 5.9, there exist $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ maximal ideals of $R$ such that $\Lambda/X=\prod_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/X_{\mathfrak{m}_{i}}$. Thus, it is a finite product of semiperfect rings. Therefore, $\Lambda/X$ is semiperfect.
@@
The next lemma will be useful for checking when the criteria of Lemma 8.1 is satisfied.
@@
which has a solution in $\mathbb{N}$ if and only if the equation $my-sx_{\ell}=r_{\ell}$ has a solution in $\mathbb{N}$. Also, it is clear that the solutions are related, as claimed in the statement.
@@
Under the Notation 7.4, the following are equivalent
@@
    For every family of natural numbers $a_{1},a_{2},\dots,a_{\ell}$, $1\leq a_{i}\leq t_{i}$, and for every $j\in\{1,\dots,\ell\}$ and $1\leq b\neq a_{j}\leq t_{j}$, there are non-negative integers $x_{1},x_{2},\dots,x_{\ell}$ such that $r_{j,b}+x_{j}r_{j,a_{j}}=x_{i}r_{i,a_{i}}$ for every $i\neq j$.
@@
    For every family of natural numbers $a_{1},a_{2},\dots,a_{\ell}$, $1\leq a_{i}\leq t_{i}$, and for every $j\in\{1,\dots,\ell\}$ and $1\leq b\neq a_{j}\leq t_{j}$, $\gcd(\mathrm{lcm}(\{r_{1,a_{1}},\dots,r_{\ell,a_{\ell}}\}\setminus\{r_{j,a_{j}}\}),r_{j,a_{j}})$ divides $r_{j,b}$.
@@
Proof. Assume $(i)$ is true, so finitely generated projective modules of $\Lambda$ modulo a (minimal) element of $\mathcal{T}(\Lambda)\setminus\{0\}$ can be lifted to finitely generated projective $\Lambda$-modules.
@@
By Remark 6.10, $\Lambda$ has an ideal $I\in\mathcal{T}(\Lambda)$ such that $I_{\mathfrak{m}_{i}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}_{i}}}(U_{i,a_{i}})$ for each $i=1,\dots,\ell$. As $\Lambda_{Q}$ is simple artinian and $R$ is $h$-local, $\Lambda/I\cong\bigoplus_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$. Therefore, every (finitely generated) projective $\Lambda/I$-module is of the form $\bigoplus_{i=1}^{\ell}M_{i}$, where $M_{i}$ is a (finitely generated) projective $(\Lambda/I)_{\mathfrak{m}_{i}}$-module. Note that $(\Lambda/I)_{\mathfrak{m}_{i}}\cong\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$, which is a semiperfect ring, so projective modules over it are described in Proposition 6.2 (iv).
@@
Consider a projective $\Lambda/I$-module $P^{\prime}$ such that $P_{\mathfrak{m}_{j}}^{\prime}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$ and $P_{\mathfrak{m}_{i}}^{\prime}=0$ if $i\neq j$. By assumption, there exists $P$ a finitely generated projective $\Lambda$-module such that $P/PI\cong P^{\prime}$. Then $P_{\mathfrak{m}_{i}}\cong U_{i,a_{i}}^{x_{i}}$ if $i\neq j$ and $P_{\mathfrak{m}_{j}}\cong U_{j,b}\oplus U_{j,a_{j}}^{x_{j}}$. Since $(P_{\mathfrak{m}_{i}})_{Q}\cong P_{Q}\cong(P_{\mathfrak{m}_{j}})_{Q}$, we obtain $x_{i}r_{i,a_{i}}=r_{j,b}+x_{j}r_{j,a_{j}}$. So $(ii)$ holds.
@@
Conversely, let $I$ be a minimal non-zero trace of $\Lambda$. Then $I_{\mathfrak{m}_{i}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}_{i}}}(U_{i,a_{i}})$ for some $a_{i}$, $i\in\{1,\dots,\ell_{i}\}$. Recall that $\Lambda/I\cong\bigoplus_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$ is a semiperfect ring. Hence, it is sufficient to lift an indecomposable projective $\Lambda/I$-module to a finitely generated projective $\Lambda$-module. So let $P^{\prime}$ be a finitely generated indecomposable projective $\Lambda/I$-module. Regarding the decomposition of $\Lambda/I$, $P^{\prime}_{\mathfrak{m}_{i}}=0$ for all $i$’s with exactly one exception. Assume $P_{\mathfrak{m}_{j}}^{\prime}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$, where $b\neq a_{j}$. Take $x_{1},\dots,x_{\ell}\in\mathbb{N}_{0}$ such that the equations in $(ii)$ are satisfied. Let $r=r_{j,b}+x_{j}r_{j,a_{j}}$.
@@
For $i=1,\dots,\ell$ set $U(\mathfrak{m}_{i})=U_{i,a_{i}}^{x_{i}}$ if $i\neq j$ and $U(\mathfrak{m}_{j})=U_{j,b}\oplus U_{j,a_{j}}^{x_{j}}$. For $\mathfrak{n}\in\operatorname{mSpec}(R)\setminus\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$, let $U(\mathfrak{n})$ be a projective $\Lambda_{\mathfrak{n}}$-module such that $U(\mathfrak{n})_{Q}$ has length $r$ (it exists by Lemma 6.9). Because of the choice of $x_{1},\dots,x_{\ell}$, we can use Proposition 7.3 to show the existence of a finitely generated projective $\Lambda$-module $P$ such that $P_{\mathfrak{m}}\cong U(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. Then $P_{\mathfrak{m}_{i}}I_{\mathfrak{m}_{i}}=P_{\mathfrak{m}_{i}}$ if $i\neq j$ and $P_{\mathfrak{m}_{j}}/P_{\mathfrak{m}_{j}}I_{\mathfrak{m}_{j}}\cong U_{j,b}/U_{j,b}I_{\mathfrak{m}_{j}}$. Therefore $P/PI\cong\bigoplus_{i=1}^{\ell}P_{\mathfrak{m}_{i}}/P_{\mathfrak{m}_{i}}I_{\mathfrak{m}_{i}}\cong P^{\prime}$. This proves $(ii)\Rightarrow(i)$.
@@
Under the assumptions in Notation 7.4, let $\ell=2$, and assume that $U_{1,1},U_{1,2}$ and $U_{2,1},U_{2,2}$ are representatives of the finitely generated indecomposable projective right modules over $\Lambda_{\mathfrak{m}_{1}}$ and $\Lambda_{\mathfrak{m}_{2}}$, respectively. Suppose that
@@
By Examples 7.5, such a locally semiperfect algebra $\Lambda$ exists and can be even taken to be a finitely generated $\mathbb{Z}$-algebra.
@@
Recall that if $\Lambda$ is taken to be finitely generated over a noetherian domain $R$ then, by Proposition 7.2, $\mathcal{A}\cong V(\Lambda)$.
@@
Under the assumptions in Notation 7.4, there are examples of algebras $\Lambda$ where there is a projective $\Lambda$-module which is not a direct sum of finitely generated modules. Set now
@@
By Examples 7.5, we know there exist locally semiperfect algebras $\Lambda$ with such invariants.
@@
Let $P]\in B(\Lambda)$ be such that $\Phi^{\prime}([P])=((0,\infty),(1,\infty))$. Then there is no $\mathbf{X}\in\mathbb{N}_{0}^{2}\times\mathbb{N}_{0}^{2}$ of the form $\mathbf{X}=((0,x_{1,2}),(1,x_{2,2}))$ being a solution of (14) since $8x_{1,2}\equiv 0\mod 4$ and $2+4x_{2,2}\equiv 2\mod 4$. Therefore, there is no finitely generated projective $\Lambda$-module $P^{\prime}$ such that $\Phi^{\prime}([P^{\prime}])=((0,x_{1,2}),(1,x_{2,2}))$, so $P$ is a countably generated projective $\Lambda$-module which is not a direct sum of finitely generated modules.
Let $P_{1}$ be a finitely generated $\Lambda$-module such that $\Phi^{\prime}([P_{1}])=((0,1),(0,2))_{8}$, and let $P_{2}$ be such that $\Phi^{\prime}([P_{2}])=((0,1),(2,1))_{8}$. By Proposition 7.1, $P\cong P\oplus P_{1}^{(\omega)}$ and $P^{2}\cong P_{2}\oplus P_{1}^{(\omega)}$. Thus $P^{2}$ is a direct sum of finitely generated projective modules.